OpenStudy (anonymous):

Express the repeating decimal as a fraction in lowest terms. 0.22=22/100+22/10,000+22/1,000,000+....

OpenStudy (anonymous):

So you want to write \(0.\overline{2}\) as a fraction, right?

OpenStudy (anonymous):

One way to do it is by writing it as a sum of fractions like you did: \[0.\overline{2}=\frac{2}{10}+\frac{2}{100}+\frac{2}{1000}+\cdots\] which is an infinite geometric series, \[\large 0.\overline{2}=\sum_{n=1}^\infty2\left(\frac{1}{10}\right)^n\] Have you worked with infinite series yet?

OpenStudy (anonymous):

Yes, i want to write it as a fraction in lowest terms

OpenStudy (anonymous):

Okay, but before I continue, do you know what infinite series are? There's another somewhat simpler method we can try if that's the case.

OpenStudy (anonymous):

is it infinity numbers?

OpenStudy (anonymous):

"Series" refers to a summation. "Infinite series" means we're adding up an infinite number of numbers. \[0.\overline{2}=\frac{2}{10}+\frac{2}{100}+\frac{2}{1000}+\frac{2}{10000}+\frac{2}{100000}+\cdots\] The \(\cdots\) means you go on and on with the pattern. The \(\sum\) form is a concise way of expressing this infinite sum. If you've never used the symbol \(\sum\) before, it sounds like you haven't learned about infinite series yet.

OpenStudy (anonymous):

yes, i have used the symbol before (sum)

OpenStudy (anonymous):

Okay, so do you know that \[\large\sum_{n=0}^\infty ar^n\] is a geometric series with common ratio \(r\)? And that \[\large\sum_{n=0}^\infty ar^n=\frac{a}{1-r}~~\text{for}~~|r|<1~~?\]

OpenStudy (anonymous):

We just started that chapter yesterday, didn't really understand it yet.

OpenStudy (anonymous):

Which part, specifically, do you not understand? I'll try to explain anything you might not get.

OpenStudy (anonymous):

how to write the equation into lowest terms

OpenStudy (anonymous):

I had not seen that formula before

OpenStudy (anonymous):

\[\sum_{i=n}^{4}1/3i\]

OpenStudy (anonymous):

is that the same, because that's all we did so far? like find the indicated sum

OpenStudy (anonymous):

That's a finite sum. It's kind of the same but there's not infinite number of terms.

OpenStudy (anonymous):

Sorry about the delayed responses. I'm a bit busy. If you have the time, feel free to watch this video. It'll explain everything I wanted to touch on: https://www.khanacademy.org/math/precalculus/seq_induction/infinite-geometric-series/v/repeating-decimal-geometric-series